### v1

@@ 3.3 Examples
Below are two examples of how the prior variance matrix $\Sigma$ encodes the structural relationships among the states. The first example generalizes the setting described in Section 2. It builds $\Sigma$ from first principles, starting with its eigenvalues and eigenvectors. The second example builds $\Sigma$ from knowledge of how the states are generated, then derives its eigenvalues and eigenvectors.
@@
The eigenvalues of (7) have mean $\sigma^{2}$, whereas the eigenvalues of (8) have mean $\nu^{2}(K+1)/2$. Choosing $\nu^{2}=2\sigma^{2}/(K+1)$ equates these two means but does not equate the eigenvalues’ distributions, nor the KL divergences $\mathcal{D}_{\text{KL}}(\mathbb{P}\parallel\mathbb{P}^{(0)})$ those distributions imply. We illustrate this fact in Figure 1. It shows that assuming states follow a random walk is equivalent, in terms of how much prior structure it imposes, to assuming a large pairwise correlation.151515 Callander, (2011) and others use Brownian motions (the continuous-time analogues of random walks) as tools for modeling “complexity.” They define “complex” environments as those in which only local learning is possible: learning a state provides some information about nearby states but little about distant states (see also Bardhi, (2024)). The limiting case is when learning a state provides no information about others; in our framework, this happens when the states are uncorrelated. Yet Figure 1 suggests that Brownian motions are as structurally restrictive as assuming states are highly correlated. This is especially true when there are many states: if $K=5$, then the equivalent correlation is about $0.82$; if $K=50$, then it is about $0.97$.
@@ 4 Preliminaries
This section contains definitions and preliminary results that we draw upon in later sections. First, we characterize the optimal action vector (3) and the posterior expected loss it induces. This loss depends on the posterior variance of the unknown vector $\theta$ given the sample $\mathcal{S}$}. We define the “value of $\mathcal{S}$}” to be the difference between the prior and posterior expected losses, and derive sharp lower and upper bounds on this value. Finally, we formalize what it means for the eigenvalues of the prior variance matrix to be “more spread out.”
4.1 Optimal actions and expected losses
Let $\operatorname{\mathbb{V}}$ take variances with respect to the prior distribution $\mathbb{P}$. Lemma 1 characterizes the optimal action vector (3) and the posterior expected loss it induces.161616 Lemma 1 holds even when $\mathbb{P}$ is not normal. We assume $\mathbb{P}$ is normal so that we can derive closed-form expressions for the posterior variances of $\theta_{1},\ldots,\theta_{K}$ and, thus, the value (11) of $\mathcal{S}$}. This vector equals the posterior mean of $\theta$. It induces a posterior expected loss equal to the mean of the posterior variances of $\theta_{1},\ldots,\theta_{K}$.
@@ 4.1 Optimal actions and expected losses
4.2 Posterior variance and Gram matrices
The expected loss (9) depends on the trace
of the posterior variance matrix $\operatorname{\mathbb{V}}(\theta\mid{\mathcal{S}})$. Lemma 2 uses the prior variance matrix $\Sigma$ and the $K\times K$ “Gram matrix”
$G\equiv\sum_{i=1}^{n}w^{(i)}(w^{(i)})^{T}$
to characterize $\operatorname{\mathbb{V}}(\theta\mid{\mathcal{S}})$.
@@ 4.2 Posterior variance and Gram matrices
4.3 Value of information
If the agent did not observe the sample $\mathcal{S}$}, then, by Lemma 1, his expected loss would equal the mean of the prior variances of $\theta_{1},\ldots,\theta_{K}$. Observing $\mathcal{S}$} lowers his minimized expected loss by
$\pi({\mathcal{S}})$ $\equiv\min_{a^{\prime}\in\mathbb{R}^{K}}\,\operatorname{\mathbb{E}}[L(\theta,a^{\prime})]-\min_{a^{\prime}\in\mathbb{R}^{K}}\,\operatorname{\mathbb{E}}[L(\theta,a^{\prime})\mid{\mathcal{S}}]$
$=\frac{1}{K}\sum_{k=1}^{K}\mathopen{}\mathclose{{\left(\operatorname{\mathbb{V}}(\theta_{k})-\operatorname{\mathbb{V}}(\theta_{k}\mid{\mathcal{S}})}}\right).$ (11)
We call $\pi({\mathcal{S}})$ the “value of $\mathcal{S}$}.”171717 Raiffa and Schlaifer, (1961, p. 90) define a similar object and call it the “(expected) value of sample information.” It is non-negative, grows as $\mathcal{S}$} grows, and shrinks as $\sigma_{u}^{2}$ grows (see Proposition A1). Intuitively, if the agent has more information, then he can take actions that estimate the states more accurately.
We can write (11) in terms of the traces of the prior and posterior variance matrices:
Thus, by Lemma 2, the value of $\mathcal{S}$} depends on the Gram matrix $G$. This matrix is symmetric and positive semi-definite. So, by the spectral theorem, there is a $K\times K$ diagonal matrix
@@ 4.4 Optimal samples
of the posterior variance matrix falls. This trace depends on the eigenvalues $\delta_{1},\ldots,\delta_{K}$ of $G$, which are non-negative, non-increasing, and sum to nn.181818 Indeed $\sum_{k=1}^{K}\delta_{k}=\operatorname{tr}(G)=\operatorname{tr}\mathopen{}\mathclose{{\left(\sum_{i=1}^{n}w^{(i)}(w^{(1)})^{T}}}\right)\overset{\star}{=}\sum_{i=1}^{n}\operatorname{tr}\mathopen{}\mathclose{{\left((w^{(i)})^{T}w^{(i)}}}\right)\overset{\star\star}{=}n,$ where $\star$ uses the linearity and cyclic property of matrix traces, and $\star\star$ uses the fact that $\lVert w^{(i)}\rVert=1$ for each $i$. So $\pi({\mathcal{S}})$ is maximized when $\delta_{1},\ldots,\delta_{K}$ solve
@@
on the value of $\mathcal{S}$}.191919 Proposition 2 echoes Liang et al.,’s (2022) Theorem 1, which says that if there are two unknown states (which Liang et al., call “attributes”), then one should prioritize learning about the state with more prior variance.
@@
for each $k\in\{1,\ldots,K\}$. Then $\pi({\mathcal{S}})\leq\pi^{*}$ with equality if
@@
We call the sample “optimal” if it induces the Gram matrix (18). The agent can construct such a sample as follows: for each $k\in\{1,\ldots,K\}$, collect $\delta_{k}^{*}$ observations with covariate $v_{k}$.202020 This may be infeasible for two reasons: (i) the eigenvalues $d_{1}^{*},\ldots,d_{K}^{*}$ may not be integers; (ii) the agent may not be able to choose $v_{1},\ldots,v_{K}$ as covariates (since, e.g., it would require him to combine negative quantities of fertilizers). We abstract from these issues for convenience and expositional clarity. Then $\mathcal{S}$} contains information about the $R^{*}$} most influential features only. This is the optimal way for the agent to “spend his data budget”: by learning about the features that matter and ignoring those that do not. The number $R^{*}$} of features that “matter” grows as his budget (i.e., the sample size $n$) grows. We call $R^{*}$} the “rank” of an optimal sample because it is the rank of the Gram matrix (18).
@@
and trace212121 The two terms on the RHS of (19) correspond to the “sampling” and “extrapolation” errors discussed in Appendix Section A2.3.
@@
The eigenvalues of $\operatorname{\mathbb{V}}(\theta\mid{\mathcal{S}})$ are the posterior variances of the unknown coefficients $\gamma_{1},\ldots,\gamma_{K}$. So if $\mathcal{S}$} is optimal, then it equates the posterior variances of $\gamma_{1},\ldots,\gamma_{R^{*}}$ to each other and the posterior variances of $\gamma_{{R^{*}}+1},\ldots,\gamma_{K}$ to their prior variances.222222 This equality of large eigenvalues and ignorance of small eigenvalues is reminiscent of Arrow,’s (1963) theorem on the optimality of deductible insurance contracts. Such contracts second-degree stochastically dominate all other contracts with the same premia (Gollier and Schlesinger,, 1996). They provide full coverage against risks above a minimum threshold. Similarly, optimal samples provide “full coverage against posterior variance” above a minimum threshold. Intuitively, the agent has a target variance and designs $\mathcal{S}$} so as to bring the posterior variances of $\gamma_{1},\ldots,\gamma_{K}$ below that target.232323 This strategy is called “reverse water-filling” in rate-distortion theory—see Cover and Thomas, (2006, Chapter 10). It also appears in Ilut and Valchev,’s (2025) model of abstract reasoning. This minimizes the trace (19) given the sample size $n$.
@@ 5 Value of conceptual knowledge
We formalize this idea as follows. Suppose the agent collects an optimal sample with value $\pi^{*}$ (see Section 4.4). Doing so relies on his conceptual knowledge of how the states relate structurally: on his ability to identify the most influential features and focus on them when collecting data. If he lacked this ability, then he would use the “naïve” prior $\mathbb{P}^{(0)}=\mathcal{N}(\mu,\Sigma^{(0)})$ described in Section 3.2. The prior variance matrix $\Sigma^{(0)}=\overline{\lambda}I_{K}$ would have equal eigenvalues $\lambda_{1}^{(0)}=\cdots=\lambda_{K}^{(0)}=\overline{\lambda}$. So, by analogy to (15) and (16), the agent’s optimal sample would have rank
@@
indexes the precision of the data relative to the agent’s prior. The difference
@@
between $\pi^{*}$ and $\pi^{(0)}$ equals the decline in the agent’s posterior expected loss from knowing how the states relate and using that knowledge to collect information optimally. Accordingly, we call $\Pi$ the “value of conceptual knowledge.” It depends on the eigenvalues $\lambda_{1},\ldots,\lambda_{K}$ of the “true” prior variance matrix $\Sigma$ and on the precision parameter $\tau$ (which jointly determine $R^{*}$}). We characterize this dependence in Theorems 1 and 2.
@@
    equals zero when $\lambda_{1},\ldots,\lambda_{K}$ are equal, and
@@
    does not fall when $\lambda_{1},\ldots,\lambda_{K}$ undergo a MPS.
@@
There is a finite threshold $\tau^{\prime}\geq 0$ such that $\Pi$ is increasing in $\tau$ if and only if $\tau<\tau^{\prime}$. This threshold equals zero if and only if $\lambda_{1},\ldots,\lambda_{K}$ are equal. Moreover,
@@
Theorem 1 says that conceptual knowledge is more valuable when states are more reducible. If a few features contribute most of states’ prior variances, then the agent gains a lot from identifying those features and collecting data on them (i.e., “asking the right questions”). In contrast, if every feature contributes equally, then he gains nothing from identifying those features because he collects the same optimal sample that he would collect if he was naïve.
Theorem 2 says that the value of conceptual knowledge is non-monotone in $\tau$, which indexes the agent’s access to information. This is because raising $\tau$ has two effects:
  1. 1.
    it gives the agent more information about the features on which he collects data, raising the gain from identifying the most influential features and collecting data on them;
  2. 2.
    it leads the agent to collect data on more features (i.e., it raises $R^{*}$}), lowering the gain from identifying the most influential features.
The first effect dominates the second precisely when $\tau<\tau^{\prime}$. The threshold $\tau^{\prime}$ equals zero if and only if $\lambda_{1},\ldots,\lambda_{K}$ are equal, in which case conceptual knowledge has no value because it does not change the optimal sample from what a naïve agent would collect.
Theorem 2 also says that the value of conceptual knowledge vanishes as the agent’s access to information grows without bound. This is because the agent’s posterior becomes less dependent on his prior as $\tau$ grows and is independent in the limit as $\tau\to\infty$. Thus, intuitively, having access to unlimited data washes out the benefit of knowing what data to collect. However, this relies on the agent having unrestricted access: he must be able to construct a sample ${\mathcal{S}}=\{(w^{(n)},y^{(n)})\}_{i=1}^{n}$ such that the covariates $w^{(1)},\ldots,w^{(n)}$ span the $K$-dimensional Euclidean space containing the state vector $\theta$. If the covariates do not span $\mathbb{R}^{K}$, then $\mathcal{S}$} may contain no information about some components of $\theta$, and so the agent’s posterior expected loss may be arbitrarily large and the value of $\mathcal{S}$} may be arbitrarily small. We illustrate this possibility in Appendix Section A2.3.
As an illustration of Theorems 1 and 2, consider the prior variance matrix (7) constructed in Example 1. Its eigenvalues are equal when $\rho=0$ and undergo a MPS when $\rho\in[0,1)$ rises. So, by Theorem 1, the value $\Pi$ of conceptual knowledge equals zero when $\rho=0$ and is non-decreasing in $\rho$. Moreover, by Theorem 2, there is a threshold $\tau^{\prime}\geq 0$ such that $\Pi$ is increasing in $\tau$ if and only if $\tau<\tau^{\prime}$. We characterize this threshold below. It equals zero when $\rho=0$ and rises when $\rho$ rises, consistent with Theorem 2.
@@
Suppose $\theta$ has prior variance (7) with $\sigma^{2}>0$ and $\rho\in[0,1)$. Then $\Pi$
@@
    is increasing in $\tau$ if and only if
    $\tau<\frac{\rho K}{1+\rho(K-1)}.$
@@
6 Deeper knowledge
Now we study the value of “deepening” the agent’s conceptual knowledge. We model this process as follows. Suppose the agent knows the trace

### v2

@@ 3.4 Examples
Below are two examples of how the prior variance matrix $\Sigma$ encodes conceptual knowledge about the states. The first example generalizes the setting described in Section 2. It builds $\Sigma$ from first principles, starting with the eigenvalues and eigenvectors. The second example builds $\Sigma$ from knowledge of how the states are generated, then derives the eigenvalues and eigenvectors.
@@
The eigenvalues of (8) have mean $\sigma^{2}$, whereas the eigenvalues of (9) have mean $\nu^{2}(K+1)/2$. Choosing $\nu^{2}=2\sigma^{2}/(K+1)$ equates these two means but does not equate the eigenvalues’ distributions, nor the KL divergences $\mathcal{D}_{\text{KL}}(\mathbb{P}\parallel\mathbb{P}^{(0)})$ those distributions imply. We illustrate this fact in Figure 1. It shows that assuming states follow a random walk is equivalent, in terms of how much prior structure it imposes, to assuming a large pairwise correlation. [^22] Callander, (2011) and others use Brownian motions (the continuous-time analogues of random walks) as tools for modeling “complexity.” They define “complex” environments as those in which only local learning is possible: learning a state provides some information about nearby states but little about distant states (see also Bardhi, (2024)). The limiting case is when learning a state provides no information about others; in our framework, this happens when the states are uncorrelated. Yet Figure 1 suggests that Brownian motions are as structurally restrictive as assuming states are highly correlated. This is especially true when there are many states: if $K=5$, then the equivalent correlation is about $0.82$; if $K=50$, then it is about $0.97$.
@@ 4 Preliminaries
This section contains preliminary results that we draw upon in later sections. We characterize the optimal action vector (3), the posterior expected loss it induces, and the value (4) of the sample $\mathcal{S}$}. We establish sharp lower and upper bounds on this value, and explain how the agent constructs “optimal” samples. Finally, we formalize what it means for the eigenvalues of the prior variance matrix to be “more spread out.”
4.1 Value of $\mathcal{S}$}
Let $\operatorname{\mathbb{V}}$ take variances with respect to the prior distribution $\mathbb{P}$. Lemma 1 characterizes the optimal action vector (3) and the posterior expected loss it induces. [^23] Lemma 1 holds even when $\mathbb{P}$ is not normal. We assume $\mathbb{P}$ is normal so that we can derive closed-form expressions for the posterior variances of $\theta_{1},\ldots,\theta_{K}$ and, thus, the value (12) of $\mathcal{S}$}. This vector equals the posterior mean of $\theta$. It induces a posterior expected loss equal to the mean of the posterior variances of $\theta_{1},\ldots,\theta_{K}$.
@@ 4.1 Value of $\mathcal{S}$}
If the sample $\mathcal{S}$} was empty, then (10) would equal the minimized prior expected loss
$\min_{a^{\prime}\in K}\operatorname{\mathbb{E}}[L(\theta,a^{\prime})]=\frac{1}{K}\sum_{k=1}^{K}\operatorname{\mathbb{V}}(\theta_{k}).$ (11)
Substituting (10) and (11) into (4) yields and expression for the sample’s value $\pi({\mathcal{S}})$ in terms of the states’ prior and posterior variances:
$\pi({\mathcal{S}})=\frac{1}{K}\sum_{k=1}^{K}\mathopen{}\mathclose{{\left(\operatorname{\mathbb{V}}(\theta_{k})-\operatorname{\mathbb{V}}(\theta_{k}\mid{\mathcal{S}})}}\right).$ (12)
Intuitively, the sample is valuable insofar as it lowers states’ variances, allowing the agent to estimate them more accurately. Moreover, since $v_{1},\ldots,v_{K}$ are known and orthonormal, we have [^24] We have $\theta=V\gamma$, where $V$ is the orthogonal matrix with known columns $v_{1},\ldots,v_{K}$ and $\gamma$ is the vector of unknown coefficients $\gamma_{1},\ldots,\gamma_{K}$. So the prior variances of $\theta_{1},\ldots,\theta_{K}$ and $\gamma_{1},\ldots,\gamma_{K}$ have equal sums: $\sum_{k=1}^{K}\operatorname{\mathbb{V}}(\theta_{k})=\operatorname{tr}\mathopen{}\mathclose{{\left(\operatorname{\mathbb{V}}(\theta)}}\right)=\operatorname{tr}\mathopen{}\mathclose{{\left(\operatorname{\mathbb{V}}(V\gamma)}}\right)=\operatorname{tr}\mathopen{}\mathclose{{\left(V\operatorname{\mathbb{V}}(\gamma)V^{T}}}\right)\overset{\star}{=}\operatorname{tr}\mathopen{}\mathclose{{\left(\operatorname{\mathbb{V}}(\gamma)}}\right)=\sum_{k=1}^{K}\operatorname{\mathbb{V}}(\gamma_{k}),$ where $\star$ uses the cyclic property of matrix traces and the orthogonality of $V$. Similarly, the posterior variances of $\theta_{1},\ldots,\theta_{K}$ and $\gamma_{1},\ldots,\gamma_{K}$ have equal sums. Substituting these sums into (12) yields (13).
$\pi({\mathcal{S}})=\frac{1}{K}\sum_{k=1}^{K}\mathopen{}\mathclose{{\left(\operatorname{\mathbb{V}}(\gamma_{k})-\operatorname{\mathbb{V}}(\gamma_{k}\mid{\mathcal{S}})}}\right).$ (13)
Thus, equivalently, the sample is valuable insofar as it lowers the variances of $\gamma_{1},\ldots,\gamma_{K}$.
4.2 Bounds on $\pi({\mathcal{S}})$
We can express (12) in terms of the traces of the prior and posterior variance matrices:
Lemma 2 characterizes $\operatorname{\mathbb{V}}(\theta\mid{\mathcal{S}})$ in terms of the prior variance matrix $\Sigma$ and “Gram matrix”
$G\equiv\sum_{i=1}^{n}w^{(i)}(w^{(i)})^{T}.$ (14)
@@ 4.2 Bounds on $\pi({\mathcal{S}})$
The Gram matrix (14) is symmetric and positive semi-definite. So, by the spectral theorem, there is a $K\times K$ diagonal matrix
@@ 4.3 Optimal samples
of the posterior variance matrix falls. This trace depends on the eigenvalues $\delta_{1},\ldots,\delta_{K}$ of $G$, which are non-negative, non-increasing, and sum to nn. [^25] Indeed $\sum_{k=1}^{K}\delta_{k}=\operatorname{tr}(G)=\operatorname{tr}\mathopen{}\mathclose{{\left(\sum_{i=1}^{n}w^{(i)}(w^{(1)})^{T}}}\right)\overset{\star}{=}\sum_{i=1}^{n}\operatorname{tr}\mathopen{}\mathclose{{\left((w^{(i)})^{T}w^{(i)}}}\right)\overset{\star\star}{=}n,$ where $\star$ uses the linearity and cyclic property of matrix traces, and $\star\star$ uses the fact that $\lVert w^{(i)}\rVert=1$ for each $i$. So $\pi({\mathcal{S}})$ is maximized when $\delta_{1},\ldots,\delta_{K}$ solve
@@
on the value of $\mathcal{S}$}. [^26] Proposition 2 echoes Liang et al.,’s (2022) Theorem 1, which says that if there are two unknown states (which Liang et al., call “attributes”), then one should prioritize learning about the state with more prior variance.
@@
for each $k\in\{1,\ldots,K\}$. Then $\pi({\mathcal{S}})\leq\pi^{*}$ with equality if $\mathcal{S}$} induces Gram matrix
@@
We call the sample “optimal” if it induces the Gram matrix (22). The agent can construct such a sample as follows: for each $k\in\{1,\ldots,K\}$, collect $\delta_{k}^{*}$ observations with covariate $v_{k}$. [^27] This may be infeasible for two reasons: (i) the eigenvalues $d_{1}^{*},\ldots,d_{K}^{*}$ may not be integers; (ii) the agent may not be able to choose $v_{1},\ldots,v_{K}$ as covariates (since, e.g., it would require him to combine negative quantities of fertilizers). We abstract from these issues for convenience and expositional clarity. Then the outcomes $y^{(1)},\ldots,y^{(n)}$ are pure signals of the coefficients $\gamma_{1},\ldots,\gamma_{K}$. For example, suppose $w^{(1)}=v_{1}$. Then, since $v_{1},\ldots,v_{K}$ are orthonormal, we have
$y^{(1)}$ $=\theta^{T}w^{(1)}+u^{(1)}$
$=\mathopen{}\mathclose{{\left(\sum_{k=1}^{K}\gamma_{k}v_{k}}}\right)^{T}v_{1}+u^{(1)}$
$=\gamma_{1}+u^{(1)}$
An optimal sample contains pure signals of $\gamma_{1},\ldots,\gamma_{R^{*}}$, the coefficients in (5) that contribute most to the states’ prior variances. However, it provides no information about $\gamma_{{R^{*}}+1},\ldots,\gamma_{K}$, the coefficients that contribute least to the states’ prior variances. Thus, the agent optimally “regularizes” by focusing on the coefficients that “matter” and ignoring those that do not. The number $R^{*}$} that “matter” grows as the sample size $n$ grows. We call $R^{*}$} the “rank” of an optimal sample because it is the rank of the Gram matrix (22).
@@
and trace [^28] The two terms on the RHS of (23) correspond to the “sampling” and “extrapolation” errors discussed in Appendix Section A2.3.
@@
The eigenvalues of $\operatorname{\mathbb{V}}(\theta\mid{\mathcal{S}})$ are the posterior variances of the unknown coefficients $\gamma_{1},\ldots,\gamma_{K}$. So if $\mathcal{S}$} is optimal, then it equates the posterior variances of $\gamma_{1},\ldots,\gamma_{R^{*}}$ to each other and the posterior variances of $\gamma_{{R^{*}}+1},\ldots,\gamma_{K}$ to their prior variances. [^29] This equality of large eigenvalues and ignorance of small eigenvalues is reminiscent of Arrow,’s (1963) theorem on the optimality of deductible insurance contracts. Such contracts second-degree stochastically dominate all other contracts with the same premia (Gollier and Schlesinger,, 1996). They provide full coverage against risks above a minimum threshold. Similarly, optimal samples provide “full coverage against posterior variance” above a minimum threshold. Intuitively, the agent has a target variance and designs $\mathcal{S}$} so as to bring the posterior variances of $\gamma_{1},\ldots,\gamma_{K}$ below that target. [^30] This strategy is called “reverse water-filling” in rate-distortion theory—see Cover and Thomas, (2006, Chapter 10). It also appears in Ilut and Valchev,’s (2025) model of abstract reasoning. This minimizes the trace (23) given the sample size $n$.
@@ 5 Value of conceptual knowledge
We formalize this idea as follows. Suppose the agent collects an optimal sample with value $\pi^{*}$ (see Section 4.3). Designing this sample relies on his conceptual knowledge: his mental model (5) of $\theta$ as a combination of concepts with different explanatory powers. If the agent did not have this knowledge, then he would use the “naïve” prior $\mathbb{P}^{(0)}=\mathcal{N}(\mu,\Sigma^{(0)})$ described in Section 3.2. He would assume $\theta$ has prior variance $\Sigma^{(0)}=\overline{\lambda}I_{K}$, a matrix with eigenvalues $\lambda_{1}^{(0)}=\cdots=\lambda_{K}^{(0)}=\overline{\lambda}$. So, by analogy to (19) and (20), the agent’s optimal sample would have rank
@@
indexes the precision of the data relative to the agent’s prior. This sample would provide equal information about each state: the optimal Gram matrix (22) would have equal eigenvalues. [^31] Indeed, if $\lambda_{1},\ldots,\lambda_{K}$ are equal (to $\overline{\lambda}$), then $\delta_{1}^{*},\ldots,\delta_{K}^{*}$ are also equal (to $n/K$). Intuitively, if the agent did not know which concepts had more explanatory power, then he would have no reason to prioritize some components of $\theta$ over others when collecting data, and so he would collect the same amount on every component.
The true and naïve prior variance matrices $\Sigma$ and $\Sigma^{(0)}$ have equal traces, and so replacing the true prior $\mathbb{P}=\mathcal{N}(\mu,\Sigma)$ with the naïve prior $\mathbb{P}^{(0)}=\mathcal{N}(\mu,\Sigma^{(0)})$ does not change the agent’s prior expected loss (see Section 4.1). But the two priors imply different optimal samples and different posterior expected losses (10). The difference
@@
between $\pi^{*}$ and $\pi^{(0)}$ equals the decline in the agent’s posterior expected loss from having conceptual knowledge and using it to design an optimal sample.
Accordingly, we call $\Pi$ the “value of conceptual knowledge.” This value depends on the eigenvalues $\lambda_{1},\ldots,\lambda_{K}$ of $\Sigma$ and the precision index $\tau$ (which jointly determine the rank $R^{*}$}). We characterize this dependence in Theorems 1 and 2.
@@
    equals zero when the eigenvalues $\lambda_{1},\ldots,\lambda_{K}$ are equal, and
@@
    does not fall when $\lambda_{1},\ldots,\lambda_{K}$ undergo a mean-preserving spread.
Theorem 1 says that conceptual knowledge is more valuable when states are more reducible. If a few common concepts explain most of states’ prior variances, then the agent gains a lot from identifying those concepts and learning about the corresponding coefficients $\gamma_{1},\ldots,\gamma_{R^{*}}$ (i.e., “asking the right questions”). In contrast, if every concept has the same explanatory power, then the agent gains nothing from identifying those concepts because he designs the same sample as he would if he was naïve.
@@
There is a finite threshold $\tau^{\prime}\geq 0$ such that the value $\Pi$ of conceptual knowledge is increasing in the precision index $\tau$ if and only if $\tau<\tau^{\prime}$. This threshold equals zero if and only if $\lambda_{1},\ldots,\lambda_{K}$ are equal. Moreover,
@@
Theorem 2 says that the value of conceptual knowledge is non-monotone in the precision index $\tau$ (and, thus, the sample size $n\equiv\tau\sigma_{u}^{2}/\overline{\lambda}$). This is because raising $\tau$ has two effects:
  1. (i)
    it gives the agent more information about the unknown coefficients $\gamma_{1},\ldots,\gamma_{R^{*}}$, raising the gain from knowing which concepts to focus on;
  2. (ii)
    it leads the agent to learn about more coefficients (i.e., it raises $R^{*}$}), lowering the gain from knowing which concepts to focus on.
The first effect dominates the second precisely when $\tau<\tau^{\prime}$. The threshold $\tau^{\prime}$ equals zero if and only if $\lambda_{1},\ldots,\lambda_{K}$ are equal, in which case conceptual knowledge has no value because it does not change the optimal sample from what a naïve agent would design.
Theorem 2 also says that the value of conceptual knowledge vanishes as $\tau$ (and thus $n$) grows without bound. This is because the agent’s posterior becomes less dependent on his prior as $\tau$ grows and is independent in the limit as $\tau\to\infty$. Intuitively, if the agent has infinite data, then he does not benefit from doing “pre-data PCA” (see Section 3.2) because he can do traditional (post-data) PCA. Having access to unlimited data washes out the benefit of knowing what data to collect. But this washout relies on having unrestricted access: the agent must be able to choose covariates $w^{(1)},\ldots,w^{(n)}$ that span the $K$-dimensional Euclidean space containing the state vector $\theta$. If the covariates do not span $\mathbb{R}^{K}$, then $\mathcal{S}$} may contain no information about some components of $\theta$, the agent’s posterior expected loss may be arbitrarily large, and the value of $\mathcal{S}$} may be arbitrarily small. We illustrate this possibility in Appendix Section A2.3.
As an illustration of Theorems 1 and 2, consider the prior variance matrix (8) constructed in Example 1. Its eigenvalues are equal when $\rho=0$ and undergo a MPS when $\rho\in[0,1)$ rises. So, by Theorem 1, the value $\Pi$ of conceptual knowledge equals zero when $\rho=0$ and is non-decreasing in $\rho$. Moreover, by Theorem 2, there is a threshold $\tau^{\prime}\geq 0$ such that $\Pi$ is increasing in the precision index $\tau$ if and only if $\tau<\tau^{\prime}$. We characterize this threshold below.
@@
Suppose the states $\theta_{1},\ldots,\theta_{K}$ have equal prior variances $\sigma^{2}>0$ and pairwise correlation $\rho\in[0,1)$. Then the value $\Pi$ of conceptual knowledge
@@
    is increasing in the precision index $\tau$ if and only if
    $\tau<\frac{\rho K}{1+\rho(K-1)}.$ (25)
@@
Figure 2: Values $\Pi$, $\pi^{*}$, and $\pi^{(0)}$ when $\theta$ has prior variance (8) and $(K,\sigma^{2},\sigma_{u}^{2})=(5,1,1)$
Figure 2 shows how $\Pi$ depends on $\rho$ and $\tau$. If $\rho=0$, then the values $\pi^{*}$ and $\pi^{(0)}$ of the optimal samples collected by agents with and without conceptual knowledge are equal, and so $\Pi=0$ for all $\tau>0$. In contrast, if $\rho>0$, then $\pi^{*}>\pi^{(0)}$ for all $\tau>0$ and hence $\Pi>0$ for all $\tau>0$. Both $\pi^{*}$ and $\pi^{(0)}$ grow as $\tau$ grows, and $\pi^{*}$ grows faster if and only if (25) holds. Thus $\Pi=\pi^{*}-\pi^{(0)}$ is increasing in $\tau$ if and only if (25) holds.
6 Deeper conceptual knowledge
Next, we study the value of “deepening” the agent’s conceptual knowledge. We model this process as follows. Suppose the agent knows the trace